OpenStudy (brooke..help00):

@whpalmer4 @chpatterson @jackthegreatest @sapphiremoon Question below----

OpenStudy (brooke..help00):

Ok, I don't know if it will work :/

OpenStudy (sshayer):

g(x)=f(x)+(3+|-2|)=f(x)+(3+2)=f(x)+5 k=5